Glavine & Co. produces a single product, each unit of which requires three direct labor hours (DLHs). Practical capacity (for setting the factory overhead application rate) is 52,000 DLHs, on an annual basis. The information below pertains to the most recent year:

Standard direct labor hours (DLHs) per unit produced 3.00
Practical capacity, in DLHs (per year) 52,000
Variable overhead efficiency variance $ 16,000 unfavorable (U)
Actual production for the year 15,000 units
Budgeted fixed manufacturing overhead $ 1,040,000
Standard direct labor wage rate $ 20.00 per DLH
Total overhead cost variance for the year $ 160,000 favorable (F)
Direct labor efficiency variance $ 32,000 unfavorable (U)

Required:

1. What was the actual number of direct labor hours (DLHs) worked during the year?

2. What was the standard variable overhead rate per DLH during the year?

3. What was the total overhead application rate per direct labor hour (DLH) during the year?

4. What was the total actual overhead cost incurred during the year?

5. What was the Production Volume Variance for the year? Was this variance favorable (F) or unfavorable (U)?

6. What was the total Overhead Spending Variance for the year? Was this variance favorable (F) or unfavorable (U)?
Answer #1

SOLUTION

1)Direct labor efficiency variance = SR [AH-SH]

32000 = 20 [AH- (15000*3)]

32000/20 = AH- 45000

AH = 1600+45000

= 46600 hours

2)Variable overhead efficiency variance =SVOR [AH-SH]

16000 = SVOR [46600-(15000*3)]

16000 =SVOR [46600-45000]

16000 =SVOR * 1600

16000/1600 =SVOR

Standard variable overhead rate per DLH = $ 10 per DLH

3)Budgeted fixed overhead rate per DLH = 1040000/52000 =$ 20per DLH

total overhead application rate per direct labor hour =Variable overhead rate+ fixed overhead rate

= 10+ 20

= $ 30 per DLH

4)Total overhead cost variance= Actual overhead- overhead absorbed

- 160000 = Actual overhead - [30* (15000*3)]

-160000 = actual overhead - 1350000

Actual overhead = -160000+1350000

= 1190000